# Public-private funding models in open source software development: A case study on scikit-learn

### v3

@@ 4.3 The muscle of industry: The role of the scikit-learn consortium
Some companies have explicitly sought to influence the technical roadmap. For example, Nvidia sponsors a full-time maintainer who is building a plug-in to enable alternative computational backends. Mike McCarty explained, “One thing that Nvidia is very interested in is scikit-learn’s effort to accelerate with GPUs. I would love to see at the end of the day scikit-learn to be accelerated natively with a very nice user experience, kind of like what we have with PyTorch, which is the canonical example of a library that you can install and if you happen to have a GPU, you can accelerate it without changing your code.” Guillaume Lemaitre, a maintainer, described receiving funding and a maintainer at the same time as “a win-win situation.” However, sponsors’ demands are not always aligned with the norms of the project. Two maintainers explained that in the past, a sponsor had requested modifications to resources that amounted, in their view, to free advertising for the company, which was not acceptable for the maintainers. While disputes of this kind have been rare, one maintainer mentioned that they can lead to the termination of their sponsorships and one company left the consortium following a disagreement.
@@
In order to safeguard community interests, the maintainers have designed governance protocols to limit the powers of the consortium. Adrin explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…The technical committee doesn’t do much in the project because the power is given to the [maintainer] team and that’s by design.” These governance measures are important to maintain the trust of the community and to mitigate fears that the project had been bought out. François explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop about MLOps since this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way; we will say, ‘Okay, we can do this, we can do that.’ But it’s a discussion and in the end the team will decide if it’s worth it or not,” he explained. Through these governance measures, the maintainers ensure that the project remains community-driven while still considering and incorporating valuable feedback from its commercial sponsors.
@@ 5.1 Implications for Research
This case study on scikit-learn illustrates the role that funding can play in supporting the sustainability and success of a non-commercial, community-led OSS project in industry-dominated fields like AI [10].
@@ 5.1.1 The design and role of public-private OSS funding models
This study makes novel contributions to the academic literature on OSS funding models. In particular, it contributes novel findings on the joint role of the public and private sector in funding the development and maintenance of an OSS project, a hitherto unexplored funding model. The analysis sheds light on the diversity of interests that public and private funders pursue in funding OSS development and maintenance, and their degree of alignment with OSS developers and maintainers. Extending prior work on incentives in OSS development, the case study illustrates that while the maintainers are primarily motivated by intrinsic incentives, such as their personal interest and values [91], companies primarily pursue economic and technological goals [58], such as maintenance and feature development, whilst governments are concerned with political and economic policy goals, including digital sovereignty [1], national competitiveness [3], and economic growth [73]. While the stakeholders pursue different goals, they are not necessarily in conflict, as they share the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not share the government’s policy goals or digital sovereignty concerns, its long-term investment in scikit-learn as a digital public good aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government do not share the companies’ business interests, the project benefits from companies’ feedback and funding for maintaining scikit-learn and expanding its functionalities.
In addition, it highlights the integral role played by the Inria Foundation as the institutional home of the scikit-learn consortium and the scikit-learn project’s governance protocols, which have served as crucial mechanisms for the maintainers to contain the influence of funders and uphold the project’s community ethos. This finding extends prior work on the role of foundations as “boundary organisations” that enable collaboration between OSS developers and for-profit companies with new evidence of the role of these institutions in handling the politics of money as an instrument of influence in community-led OSS projects [50, 24]. Moreover, the diversified funding model not only ensures financial stability for the project but also acts as a safeguard when disagreements arise between maintainers and funders. This diversification protects the project from the risk of over-reliance on the support of a single funder.
@@ 5.1.2 Research design for future research on OSS funding
There is a dearth of research on OSS funding, relative to other topics in the extensive literature on the economics of OSS [92]. With the emergence of new funding models, from FOSS Contributor Funds to governmental funds like the STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s), amongst others. To this end, this paper contributes novel insights and a research design that can inform future studies on OSS funding. As discussed below, while the uniqueness of scikit-learn limits the generalisability of the case-specific findings, there are aspects of the scikit-learn experience that will resonate with OSS developers in general, such as the processes of balancing the varied interests of funders while preserving community values.
@@ 5.2.1 Implications for OSS developer communities
The experience of scikit-learn shows that a diversified funding model can act as a useful mechanism for safeguarding the project’s independence and community-driven ethos, whilst benefiting from the financial support of public and private funding, which have their relative merits. On the one hand, companies focus on specific technical problems related to maintenance and the development of new functionalities. As one maintainer explained, as users of scikit-learn, they need it to work. Maintainers shared their appreciation for the exchanges with commercial sponsors, which are “rooted in something real.” However, in some cases, companies make requests that the maintainers consider to be at odds with the interests of the project and refuse to implement, which can create tension and uncertainty over whether the company will sponsor the project in the subsequent year. On the other hand, the government’s focus on policy goals meant that initially some requests of governmental sponsors were disconnected from the needs and expertise of the maintainers. For example, initially the French government overlooked maintenance and asked the maintainers to work on DL, which was beyond their expertise. However, through an ongoing dialogue, these concerns were taken into consideration by the government’s AI strategy coordinators and the development priorities of the government were in turn adapted. Furthermore, maintainers appreciated the government’s requirement to secure additional public and private funding, which ensured the continued diversification of the funding base at a time when the French government was investing as much as €32 million into scikit-learn. As François remarked, “Overall, it’s good to have a panaché.”
@@ 5.2.2 Implications for companies
For companies, the findings are a timely reminder that by sponsoring a developer and/or funding a project, they can make a substantial difference to OSS maintainers, who struggle with limited resources and towering workloads [29]. As shown in Table 2, the scikit-learn project’s community includes several technology giants and industry leaders, such as Google, Meta, and AWS, which have not funded the project or sponsored maintainers to date. As one interviewee explained, while such companies contribute useful issues and pull requests, ultimately their contributions create more work for the already overstretched maintainers. It would be more helpful if companies such as these Big Technology companies sponsored a maintainer or funded the project by joining the consortium, which might be a low-effort intervention for a well-resourced industry giant with a high reward for the maintainers of scikit-learn.
@@ 5.2.3 Implications for governments
For governments, the findings underscore the critical importance of funding the maintenance of existing OSS and refraining from the novelty bias of exclusively funding the development of new OSS libraries or features.. This finding extends prior recommendations for the funding of scientific OSS [70]. Furthermore, it highlights the value of engaging in multi-stakeholder dialogue with OSS developers in order to inform the effective design and delivery of grants that strike a balance between the government’s policy goals and the daily realities of OSS developers. The scikit-learn maintainers explained that while they appreciated the recognition of the importance of scikit-learn in AI R&D in the French government’s AI strategy, they admitted they had faced difficulties in translating the government’s policy goals into practice. While they could understand the government’s rationale to strengthen digital sovereignty and national competitiveness, the government initially made demands that the maintainers viewd as either difficult to implement in practice or disconnected from their expertise, such as working on DL when their expertise was in ML. The maintainers suggested that this was in part due to a non-expert understanding of policymakers, which was remedied through ongoing multi-stakeholder dialogue between government officials, the scikit-learn maintainers, and other parties.
@@ 5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefitting from the feedback from domain experts as well as peer review. The combination of these steps enhanced the dependability of the study.
@@ 6 Conclusion
Funding plays a critical role in supporting the sustainability of community-led OSS projects. This case study on scikit-learn makes two key contributions with implications for research and practice. First, the study contributes novel findings on emerging OSS funding models to the academic literature. It illustrates how the maintainers of scikit-learn have weaved public and private sources of funding into their project in the interest of ensuring the continued provision of scikit-learn as a digital public good, whilst employing governance protocols to safeguard the community ethos of the project. Second, it provides practical recommendations for OSS practitioners. For OSS developer communities, it shows that the diversification of funding can act as a useful mechanism for sustaining a community-led OSS project, whilst safeguarding the community ethos of the project and its independence from the overreaching influence of single funders, whether they are a company, a government body, or other. For companies, it is a timely reminder that by sponsoring an OSS developer or funding an OSS project they can make a significant difference to maintainers. For governments, it underscores the critical importance of refraining from the novelty bias of only funding new innovation and funding the maintenance of existing OSS. Overall, this study underscores the effectiveness of diversified funding models as a pragmatic approach to sustaining community-led OSS projects in competitive industries. Ultimately, the future of OSS funding depends on the development of sustainable strategies that align the interests of funders with the needs and norms of the global developer communities that build and maintain our open source digital infrastructure.
@@ Acknowledgements
This study was supported by the UK’s Economic and Social Research Council [Grant number: ES/P000649/1].

### v4

@@ 4.3 Private Sponsorship: The Role of the scikit-learn Consortium
In practice, this dynamic has yielded a mix of synergies and tensions. In some cases, companies’ business interests are compatible with scikit-learn’s technical roadmap. For example, in addition to its consortium membership, Nvidia has sponsored a maintainer to build a plug-in that enables alternative computational backends for scikit-learn. Mike McCarty explained, “One thing that Nvidia is very interested in is scikit-learn’s effort to accelerate with GPUs. I would love to see at the end of the day scikit-learn to be accelerated natively with a very nice user experience, kind of like what we have with PyTorch, which is the canonical example of a library that you can install and if you happen to have a GPU, you can accelerate it without changing your code”. Guillaume Lemaitre, a maintainer, described Nvidia’s simultaneous sponsorship of a maintainer and consortium membership as “a win-win situation”.
In other cases, however, sponsor requests to shape the project’s direction or burnish their own reputations have run afoul of project norms. Two anonymous maintainers explained that in the past, a sponsor had requested modifications to the technical documentation that amounted, in their view, to free advertising for the company. This self-promotional request, they explained, was not acceptable, especially not in the technical documentation of the project. While disputes of this kind have been rare, one anonymous maintainer mentioned that disputes had led to the termination of sponsorship by one company in the past. To manage the interests and demands of consortium members, the scikit-learn maintainers employ carefully designed governance protocols that limit funders’ influence and prioritise community control. The Technical Committee, for instance, is barred from overriding decisions that achieve rough consensus amongst the maintainers. Adrin Jalali explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…because the power is given to the [maintainers] and that’s by design”.
@@
These governance measures are important to maintain the trust of the community and to mitigate fears that the project had been bought out. François Goupil explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop on MLOps because this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way…but it’s a discussion and in the end the team will decide if it’s worth it or not”, he explained. Through these governance protocols, the maintainers ensure that the project remains community-driven while still considering and incorporating feedback from the consortium. As François Goupil put it, “We are a community with different interests”, and preserving that pluralism is key to the project’s long-term sustainability and independence.
Finally, even if relationships with their sponsors can be troublesome at times, the maintainers were grateful for their support and for not free-riding on scikit-learn, which many companies do. During an on-site visit, François Goupil showed me a dashboard, comprising world maps of contributors and charts of contributions from companies. Upon observing that the most active organisations were Chinese and American companies (see Table 2), I asked if these insights had influenced his opinion about whether they contributed sufficiently. He responded frankly: while such companies contribute useful issues and pull requests, their contributions ultimately create more work for the maintainers. He explained that it would be more helpful if these companies sponsored a maintainer or joined the consortium in order to provide more capacity to the maintainers to process the extra workload created by these contributions.
@@ 5.1.1 The design and role of public-private OSS funding models
This study contributes novel findings on the roles of public and private funding in OSS development and maintenance. It reveals the diverse interests of funders, from the economic and technological goals of companies [48] to the policy objectives like digital sovereignty [1], national competitive in science and innovation [3], and economic growth [72] for governments. While the various stakeholders pursue different goals, they are not necessarily in conflict, as they share the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not necessarily share the government’s policy goals or digital sovereignty concerns, its long-term investment in scikit-learn aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government officials do not necessarily share the companies’ business interests in scikit-learn, the project benefits from companies’ feedback and funding for development and maintenance. The study also highlights the role of the governance mechanisms to balance the interests of the community and the companies in the consortium, extending prior work on the role of foundations as “boundary organisations” that mitigate the dominance of companies and enable collaboration between OSS developers and companies [16, 54].
@@ 5.1.2 Research design for future research on OSS funding
There is still a dearth of research on OSS funding, relative to other topics on the economics of OSS [91]. With the emergence of new funding models, from FOSS Contributor Funds to governmental funds like Germany’s STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s), amongst others. To this end, this paper contributes novel insights and a research design that can inform future studies on OSS funding. As discussed below, while the uniqueness of scikit-learn limits the generalisability of the case-specific findings, there are aspects of the scikit-learn experience that will resonate with OSS developers in general, such as the processes of balancing the varied interests of funders while preserving community values.
@@ 5.2.1 Implications for OSS Developer Communities
The case of scikit-learn shows that a diversified funding model can be an effective strategy for safeguarding a project’s independence and community-driven ethos whilst benefiting from both public and private support. Furthermore, public and private funding come with their own challenges. On the one hand, commercial sponsors often focus on specific technical issues related to maintenance and feature development, and their demands may sometimes conflict with project goals or norms. On the other hand, government funders may prioritise high-level policy objectives over the activities or expertise of the project. Navigating these tensions requires ongoing dialogue and negotiation to align goals and ensure a sustainable balance of funding sources. Ultimately, maintainers must be prepared to advocate for and defend their project’s needs and values. As the scikit-learn case demonstrates, a diverse funding portfolio can provide both stability and leverage in these discussions.
@@ 5.2.2 Implications for Companies
For companies, the findings serve as a timely reminder that sponsoring OSS maintainers or projects can make a substantial difference to maintainers, who often struggle with limited capacity and towering workloads [27]. As shown in Table 2, several major technology companies, such as Google, Meta, and AWS, have benefited significantly from scikit-learn but have not yet directly funded the project or its maintainers. While these companies may contribute useful issues and pull requests, their engagement can also create additional work for already overstretched maintainers. Sponsoring a maintainer or joining a project’s funding consortium, even at a relatively low level, could be a high-impact way for well-resourced companies to support the OSS projects that they use and depend on. By stepping up to provide more direct financial support, companies can help ensure the sustainability and health of critical OSS infrastructure.
@@ 5.2.3 Implications for Governments
For governments, the findings highlight the importance of funding the maintenance of existing, widely-used OSS infrastructure and not just the development of new libraries or features, extending prior recommendations for the funding of scientific OSS [68]. To ensure that funding is aligned with project needs, governments should engage in multi-stakeholder dialogue with OSS developers. This consultation is crucial for designing grants that balance policy goals with the expertise of maintainers. In the case of scikit-learn, initial government requests around expanding into deep learning were misaligned with the project’s scope and capabilities. Ongoing conversations were necessary to bridge this gap and ensure that funding supported both overarching policy objectives and essential project functions. Governments should proactively seek input from the OSS community to avoid such misalignments and to craft funding programs that effectively support the sustainability of critical (scientific) OSS projects.
@@ 5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefiting from the feedback from domain experts as well as peer review. The combination of these steps enhanced the dependability of the study.
@@ 6 Conclusion
This study on the funding model of scikit-learn makes two key contributions to research and practice. First, it contributes novel findings on emerging OSS funding models, illustrating how the maintainers of an impactful community-led OSS project have combined public and private funding to sustain the development and maintenance of the project, whilst safeguarding its community ethos. Second, it offers practical lessons for the funding of community-led OSS projects, and it makes recommendations to various OSS practitioners. In particular, it highlights effectiveness of diversified funding models as a method to sustain community-led OSS projects in industry-dominated R&D fields, such as AI.
@@ Acknowledgements
This study was funded by the UK’s Economic and Social Research Council [Grant number: ES/P000649/1]. The author thanks the scikit-learn maintainers and funders for participating in this study.